OpenStudy (anonymous):

Any drops a ball from an initial height of 30 feet above the ground.After how many seconds will the ball reach the ground? (Round your answer to the nearest tenth of a second.)

OpenStudy (gabby1214):

-16t^2 + 30 = 0 => 16t^2 = 30 => t^2 = 30/16 => t^2 = 15/8 t = sqrt(15/8) = 1.369 = 1.4 1.4 seconds I'm pretty sure its right but someone could correct me if i'm wrong